# Gradient estimates of the Finslerian Allen-Cahn equation

### v1

@@ 4 Global gradient estimates on compact Finsler manifolds
$\begin{split}g_{\nabla u}(\nabla^{\nabla u}H,\nabla^{\nabla u}\log w)=2&\left(\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}}\right.\\ &\quad\left.-\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}+\frac{\beta}{q}w^{-\frac{2}{q}}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}\right).\end{split}$ (4.18)
It is deduced from (4.15) by combining (4.16)-(4.18) and employing the definition of weighted Ricci curvature that
$\begin{split}\Delta^{\nabla u}H\geq&\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{(\Delta^{\nabla u}w)^{2}}{w^{2}}-2(\frac{1}{\epsilon}-1)\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+\frac{2}{q}g_{\nabla u}(\nabla^{\nabla u}H,\nabla^{\nabla u}\log w)+2\frac{Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\\ &+(4-6\frac{\beta}{q^{2}})\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}+2\beta w^{-\frac{2}{q}}(1-w^{-\frac{2}{q}}),\end{split}$ (4.19)
on $M_{u}$. Noticing from (4.1) and the expression of $H$, we get that
$\frac{\Delta^{\nabla u}w}{w}=\frac{q}{\beta}H+(\frac{q+1}{q}-\frac{q}{\beta})\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}.$ (4.20)
Plugging (4.20) into (4.19), and setting $\beta=sq^{2}$, we arrive at the following lemma.
Lemma 4.2.
Let $(M,F,\mu)$ be a forward complete Finsler metric measure manifold, and denote $M_{u}=\{x\in M\mid\nabla u(x)\neq 0\}$. For $H=\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}+\beta(1-w^{-\frac{2}{q}})$ with $\beta=sq^{2}$, it holds on $M_{u}$ that
$\begin{split}\Delta^{\nabla u}H\geq&\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{H^{2}}{s^{2}q^{2}}+\left[\frac{2(1-\epsilon)(sq+s-1)^{2}}{(N-\epsilon(N-n))s^{2}q^{2}}-2(\frac{1}{\epsilon}-1)\right]\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+\frac{4(1-\epsilon)(sq+s-1)}{((1-\epsilon)N+\epsilon n)s^{2}q^{2}}H\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}+\frac{2}{q}g_{\nabla u}(\nabla^{\nabla u}H,\nabla^{\nabla u}\log w)\\ &+2w^{-\frac{2}{q}}H+(2-6s)\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}+2\frac{Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}.\end{split}$ (4.21)
With the assistance of Lemma 4.2, we can prove the global gradient estimates on compact Finsler metric measure manifolds directly, with the weighted Ricci curvature bounded from below. Precisely, we have
Theorem 4.3.
Let $(M,F,\mu)$ be a compact Finsler metric measure manifold whose weighted Ricci curvature satisfies $Ric^{N}\geq-K$, with some positive constant $K$. Assume the bound of the reversibility on $M$ is $\rho_{0}$. Suppose $u$ is a bounded positive solution to the Finslerian Allen-Cahn equation (1.3) on $M$ with $u\leq C$ where $C$ is a positive constant. (1) If $C\leq 1$, then we have
@@
for any $10<\epsilon<$ 1. (2) If $C\geq 1$, then we have
@@
Lemma 4.2 implies that (4.21) holds in the distributional sense on $M$, i.e., for any nonnegative test function $\phi$, $H$ satisfies that
@@
where $\beta$ denotes the RHS of (4.21). Suppose $x_{0}$ is the maximum point of $H$ on $M$. Without loss of generality, one could assume $H(x_{0})\geq 0$, otherwise the result will be satisfied trivially. We claim that $\beta(x_{0})\leq 0$.
If not, $\beta$ is strictly positive at $x_{0}$, so that $\beta(x)$ is positive in a small neighborhood of $x_{0}$ on $M$, which may be denoted by $U$. Chosen a test function $\phi$ whose compact support set is contained in $U$, we know from (4.22) that $H$ is a weak, local subharmonic function in a neighborhood $V\subset U$. It is a contradiction because $x_{0}$ is a inner point of $V$. Since $\beta(x_{0})\leq 0$ and $\nabla^{\nabla u}H(x_{0})=0$, we arrive at
$\begin{split}0\geq&\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{H^{2}}{s^{2}q^{2}}+\left[\frac{2(1-\epsilon)(sq+s-1)^{2}}{(N-\epsilon(N-n))s^{2}q^{2}}-2(\frac{1}{\epsilon}-1)\right]\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+\frac{4(1-\epsilon)(sq+s-1)}{((1-\epsilon)N+\epsilon n)s^{2}q^{2}}H\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}+2w^{-\frac{2}{q}}H\\ &+(2-6s)\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}+2\frac{Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}.\end{split}$ (4.23)
We consider (4.23) in two cases: (1) $C\leq 1$ and (2) $C>1$.
@@
$\frac{Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\geq-K\frac{F^{2}(\nabla^{\nabla u}w)}{w^{2}}\geq-K\rho_{0}^{2}H.$ (4.24)
@@
$\begin{split}&2w^{-\frac{2}{q}}H+(2-6s)\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}\\ =&(4-6s)w^{-\frac{2}{q}}H+(6s-2)sq^{2}(1-w^{-\frac{2}{q}})w^{-\frac{2}{q}}\\ \geq&-(6s-4)H,\end{split}$ (4.25)
@@
We may choose $s=\frac{2}{3}$ so that the RHS of (4.25) is equal to 0, and choose $0<q\leq\frac{1}{2}\left(1+\sqrt{\frac{[N-\epsilon(N-n)]}{\epsilon}}\right)^{-1}$ so that the second term on the RHS of (4.23) in this case is nonnegative. Furthermore, the first and third terms on the RHS of (4.23) could be bounded from below by
@@
Plugging (4.24) and (4.25) into (4.23) with the above estimates gives that
@@
$H\leq\frac{4\rho_{0}^{2}Kq^{2}[N-\epsilon(N-n)]}{3(1-\epsilon)}.$ (4.26)
@@
$\begin{split}&\quad\frac{2Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\geq-2\rho_{0}^{2}K\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}\\ &\geq-\frac{2(1-\epsilon)(s-1)^{2}}{(N-\epsilon(N-n))s^{2}q^{2}}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}-\frac{(N-\epsilon(N-n))s^{2}q^{2}}{2(1-\epsilon)(s-1)^{2}}\rho_{0}^{4}K^{2},\end{split}$ (4.27)
@@
$(6s-2)\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}\leq(3s-1)^{2}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}+C^{4},$ (4.28)
@@
$\frac{2(1-\epsilon)}{(N-\epsilon(N-n))}\frac{s-1}{sq}\geq\frac{1}{\epsilon}-1+\frac{(3s-1)^{2}}{2}.$ (4.29)
We plug (4.27)-(4.29) into (4.23) to obtain that
@@
$H\leq\frac{[N-\epsilon(N-n)]}{2(1-\epsilon)(s-1)}s^{2}q^{2}\rho_{0}^{2}K+\sqrt{\frac{[N-\epsilon(N-n)]}{2(1-\epsilon)}}sqC^{2}.$ (4.30)
@@
Remark 4.4.
Theorem 1.1 follows by taking $\epsilon=\frac{1}{2}$ and $q=\frac{1}{\sqrt{N+n}}$ in case 1. as well as taking $s=2$, $\epsilon=\frac{1}{2}$ and $q=\frac{2}{27(N+n)}$ in case 2. of Theorem 4.3.
@@ 5 Local gradient estimates on forward complete Finsler manifolds
In this section, we prove the local gradient estimates on forward complete Finsler metric measure manifolds with the assistance of Lemma 4.2 and the Comparison theorem.
Theorem 5.1.
Let $(M,F,\mu)$ be a complete noncompact Finsler metric measure manifold. Denote by $B(p,2R)$ the forward geodesic ball of radius $2R$ centered at $p$ on $M$. Suppose the mixed weighted Ricci curvature satisfies that ${}^{m}Ric^{N}\geq-K(2R)$ in $B(p,2R)$ with $K(2R)\geq 0$, and the misaligment $\alpha$ satisfies $\alpha\leq A(2R)$ in $B(p,2R)$. Moreover, the non-Riemannian tensors satisfy $F(U)+F^{*}(\mathcal{T})+F(\mathrm{div}C(V))\leq K_{0}$. Let $u$ be a bounded positive solution to the Finslerian Allen-Cahn equation (1.3) in $B(p,2R)$ with $u\leq C$ where $C$ is a positive constant. (1) If $C\leq 1$, then we have
@@
in $B(p,R)$ for any $10<\epsilon<$ 1, where $C_{1},C_{2}$ are positive constants. (2) If $C>1$, then we have
@@
Let $\tilde{\phi}\in C^{2}[0,+\infty)]$ such that $\tilde{\phi}(r)=1$ when $r\leq 1$ and $\tilde{\phi}(r)=0$ whenever $r\geq 2$. Hence $0\leq\tilde{\phi}(r)\leq 1$. Moreover, we may assume that
@@
$\Delta^{\nabla u}(\phi H)=H\Delta^{\nabla u}\phi+\phi\Delta^{\nabla u}H+\frac{2}{\phi}g_{\nabla u}(\nabla^{\nabla u}(\phi H),\nabla^{\nabla u}\phi)-\frac{2H}{\phi}F_{\nabla u}^{2}(\nabla^{\nabla u}\phi),$ (5.1)
where $\nabla^{\nabla u}\phi=\tilde{\phi}^{\prime}\frac{\nabla^{\nabla u}r}{R}$, $\Delta^{\nabla u}\phi=\tilde{\phi}^{\prime\prime}\frac{F^{2}_{\nabla u}(\nabla^{\nabla u}r)}{R^{2}}+\tilde{\phi}^{\prime}\frac{\Delta^{\nabla u}r}{R}$. Plugging these two terms into (5.1) gives that
$\begin{split}\phi\Delta^{\nabla u}(\phi H)=&\phi H(\Delta^{\nabla u}\phi-\frac{2}{\phi}F_{\nabla u}^{2}(\nabla^{\nabla u}\phi))\\ &+2\phi g_{\nabla u}(\nabla^{\nabla u}(\phi H),\nabla^{\nabla u}\log\phi)+\phi^{2}\Delta^{\nabla u}H\\ \geq&\phi H\left[-\frac{\alpha C_{2}}{R^{2}}-\frac{2\alpha C_{1}^{2}}{R^{2}}-\frac{C_{1}}{R}(C(N,\alpha)\mathfrak{ct}_{l}(r)+\sqrt{\alpha}K_{0})\right]\\ &+2\phi g_{\nabla u}(\nabla^{\nabla u}(\phi H),\nabla^{\nabla u}\log\phi)+\phi^{2}\Delta^{\nabla u}H,\end{split}$ (5.2)
on $M_{u}$. (5.2) provides by taking the estimation of $\Delta^{\nabla u}H$ in Lemma 4.2 that
$\begin{split}\phi\Delta^{\nabla u}(\phi H)\geq&2\phi g_{\nabla u}(\nabla^{\nabla u}(\phi H),\nabla^{\nabla u}\log\phi)+\frac{2\phi}{q}g_{\nabla u}(\nabla^{\nabla u}(\phi H),\nabla^{\nabla u}\log w)\\ &-B\phi H+\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{(\phi H)^{2}}{s^{2}q^{2}}+2\phi^{2}\frac{Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\\ &+\phi^{2}\left[\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{(sq+s-1)^{2}}{s^{2}q^{2}}-2(\frac{1}{\epsilon}-1)\right]\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+\phi\frac{4(1-\epsilon)(sq+s-1)}{N-\epsilon(N-n)s^{2}q^{2}}(\phi H)\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}+2w^{-\frac{2}{q}}\phi^{2}H\\ &+(2-6s)\phi^{2}w^{-\frac{2}{q}}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}-\frac{2}{q}(\phi H)g_{\nabla u}(\nabla^{\nabla u}\phi,\nabla^{\nabla u}\log w),\end{split}$ (5.3)
where $B=\frac{2\alpha C_{1}^{2}}{R^{2}}+\frac{C_{1}}{R}\left[\sqrt{K(2R)C(N,A)}\coth\left(R\sqrt{\frac{K(2R)}{C(N,A)}}\right)+C_{0}(K_{0},A)\right]+\frac{\alpha C_{2}}{R^{2}}$. It could be deduced from (5.3) that
@@
$\begin{split}B\phi H\geq&\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{(\phi H)^{2}}{s^{2}q^{2}}+2\phi^{2}\frac{Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\\ &+\phi^{2}\left[\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{(sq+s-1)^{2}}{s^{2}q^{2}}-2(\frac{1}{\epsilon}-1)\right]\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+\phi\frac{4(1-\epsilon)(sq+s-1)}{N-\epsilon(N-n)s^{2}q^{2}}(\phi H)\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}\\ &+(2-6s)\phi^{2}w^{-\frac{2}{q}}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}-\frac{2}{q}(\phi H)g_{\nabla u}(\nabla^{\nabla u}\phi,\nabla^{\nabla u}\log w),\end{split}$ (5.5)
is valid at $x_{1}$. Next we also consider (5.5) in two cases: (1) $C\leq 1$ and (2) $C>1$.
@@
$\frac{2Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\phi^{2}\geq-2K(2R)\phi^{2}\frac{F^{2}(\nabla^{\nabla u}w)}{w^{2}}\geq-2AK(2R)\phi H,$ (5.6)

### v2

@@ 4 Global gradient estimates on compact Finsler manifolds
$\begin{split}g_{\nabla u}(\nabla^{\nabla u}H,\nabla^{\nabla u}\log w)=2&\left(\frac{g_{\nabla u}((\nabla^{\nabla u})^{2}w,\nabla^{\nabla u}w\otimes\nabla^{\nabla u}w)}{w^{3}}\right.\\ &\quad\left.-\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}+\frac{\beta}{q}w^{-\frac{2}{q}}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}\right).\end{split}$ (50)
It is deduced from (47) by combining (48)-(50) and employing the definition of weighted Ricci curvature that
$\begin{split}\Delta^{\nabla u}H\geq&\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{(\Delta^{\nabla u}w)^{2}}{w^{2}}-2(\frac{1}{\epsilon}-1)\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+\frac{2}{q}g_{\nabla u}(\nabla^{\nabla u}H,\nabla^{\nabla u}\log w)+2\frac{Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\\ &+(4-6\frac{\beta}{q^{2}})\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}+2\beta w^{-\frac{2}{q}}(1-w^{-\frac{2}{q}}),\end{split}$ (51)
on $M_{u}$. Noticing from (33) and the expression of $H$, we get that
$\frac{\Delta^{\nabla u}w}{w}=\frac{q}{\beta}H+(\frac{q+1}{q}-\frac{q}{\beta})\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}.$ (52)
Plugging (52) into (51), and setting $\beta=sq^{2}$, we arrive at the following lemma.
Lemma 9.
Let $(M,F,\mu)$ be a forward complete Finsler metric measure space, and denote $M_{u}=\{x\in M\mid\nabla u(x)\neq 0\}$. For $H=\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}+\beta(1-w^{-\frac{2}{q}})$ with $\beta=sq^{2}$, it holds on $M_{u}$ that
$\begin{split}\Delta^{\nabla u}H\geq&\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{H^{2}}{s^{2}q^{2}}+\left[\frac{2(1-\epsilon)(sq+s-1)^{2}}{(N-\epsilon(N-n))s^{2}q^{2}}-2(\frac{1}{\epsilon}-1)\right]\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+\frac{4(1-\epsilon)(sq+s-1)}{((1-\epsilon)N+\epsilon n)s^{2}q^{2}}H\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}+\frac{2}{q}g_{\nabla u}(\nabla^{\nabla u}H,\nabla^{\nabla u}\log w)\\ &+2w^{-\frac{2}{q}}H+(2-6s)\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}+2\frac{Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}.\end{split}$ (53)
With the assistance of Lemma 9, we can prove the global gradient estimates on compact Finsler metric measure spaces directly, with the weighted Ricci curvature bounded from below. Precisely, we have
Theorem 10.
Let $(M,F,\mu)$ be a compact Finsler metric measure space whose weighted Ricci curvature satisfies $Ric^{N}\geq-K$, with some positive constant $K$. Assume the bound of the reversibility on $M$ is $\rho_{0}$. Suppose $u$ is a bounded positive solution to the Finslerian Allen-Cahn equation (3) on $M$ with $u\leq C$ where $C$ is a positive constant.
(1) If $C\leq 1$, then we have
@@
for any $10<\epsilon<$ 1.
(2) If $C\geq 1$, then we have
@@
Lemma 9 implies that (53) holds in the distributional sense on $M$, i.e., for any nonnegative test function $\phi$, $H$ satisfies that
@@
where $\beta$ denotes the RHS of (53).
Suppose $x_{0}$ is the maximum point of $H$ on $M$. Without loss of generality, one could assume $H(x_{0})\geq 0$, otherwise the result will be satisfied trivially. We claim that $\beta(x_{0})\leq 0$.
If not, $\beta$ is strictly positive at $x_{0}$, so that $\beta(x)$ is positive in a small neighborhood of $x_{0}$ on $M$, which may be denoted by $U$. Chosen a test function $\phi$ whose compact support set is contained in $U$, we know from (54) that $H$ is a weak, local subharmonic function in a neighborhood $V\subset U$. It is a contradiction because $x_{0}$ is a inner point of $V$.
Since $\beta(x_{0})\leq 0$ and $\nabla^{\nabla u}H(x_{0})=0$, we arrive at
$\begin{split}0\geq&\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{H^{2}}{s^{2}q^{2}}+\left[\frac{2(1-\epsilon)(sq+s-1)^{2}}{(N-\epsilon(N-n))s^{2}q^{2}}-2(\frac{1}{\epsilon}-1)\right]\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+\frac{4(1-\epsilon)(sq+s-1)}{((1-\epsilon)N+\epsilon n)s^{2}q^{2}}H\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}+2w^{-\frac{2}{q}}H\\ &+(2-6s)\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}+2\frac{Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}.\end{split}$ (55)
We consider (55) in two cases: (1) $C\leq 1$ and (2) $C>1$.
@@
$\frac{Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\geq-K\frac{F^{2}(\nabla^{\nabla u}w)}{w^{2}}\geq-K\rho_{0}^{2}H.$ (56)
@@
$\begin{split}&2w^{-\frac{2}{q}}H+(2-6s)\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}\\ =&(4-6s)w^{-\frac{2}{q}}H+(6s-2)sq^{2}(1-w^{-\frac{2}{q}})w^{-\frac{2}{q}}\\ \geq&-(6s-4)H,\end{split}$ (57)
@@
We may choose $s=\frac{2}{3}$ so that the RHS of (57) is equal to 0, and choose $0<q\leq\frac{1}{2}\left(1+\sqrt{\frac{[N-\epsilon(N-n)]}{\epsilon}}\right)^{-1}$ so that the second term on the RHS of (55) in this case is nonnegative. Furthermore, the first and third terms on the RHS of (55) could be bounded from below by
@@
Plugging (56) and (57) into (55) with the above estimates gives that
@@
$H\leq\frac{4\rho_{0}^{2}Kq^{2}[N-\epsilon(N-n)]}{3(1-\epsilon)}.$ (58)
@@
$\begin{split}&\quad\frac{2Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\geq-2\rho_{0}^{2}K\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}\\ &\geq-\frac{2(1-\epsilon)(s-1)^{2}}{(N-\epsilon(N-n))s^{2}q^{2}}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}-\frac{(N-\epsilon(N-n))s^{2}q^{2}}{2(1-\epsilon)(s-1)^{2}}\rho_{0}^{4}K^{2},\end{split}$ (59)
@@
$(6s-2)\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}w^{-\frac{2}{q}}\leq(3s-1)^{2}\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}+C^{4},$ (60)
@@
$\frac{2(1-\epsilon)}{(N-\epsilon(N-n))}\frac{s-1}{sq}\geq\frac{1}{\epsilon}-1+\frac{(3s-1)^{2}}{2}.$ (61)
We plug (59)-(61) into (55) to obtain that
@@
$H\leq\frac{[N-\epsilon(N-n)]}{2(1-\epsilon)(s-1)}s^{2}q^{2}\rho_{0}^{2}K+\sqrt{\frac{[N-\epsilon(N-n)]}{2(1-\epsilon)}}sqC^{2}.$ (62)
@@
Remark 3.
Theorem 1 follows by taking $\epsilon=\frac{1}{2}$ and $q=\frac{1}{\sqrt{N+n}}$ in case 1. as well as taking $s=2$, $\epsilon=\frac{1}{2}$ and $q=\frac{2}{27(N+n)}$ in case 2. of Theorem 10.
@@ 5 Local gradient estimates on forward complete Finsler manifolds
In this section, we prove the local gradient estimates on forward complete Finsler metric measure spaces with the assistance of Lemma 9 and the Comparison theorem.
Theorem 11.
Let $(M,F,\mu)$ be a complete noncompact Finsler metric measure space. Denote by $B(p,2R)$ the forward geodesic ball of radius $2R$ centered at $p$ on $M$. Suppose the mixed weighted Ricci curvature satisfies that ${}^{m}Ric^{N}_{\nabla r}\geq-K(2R)$ in $B(p,2R)$ with $K(2R)\geq 0$, and the misaligment $\alpha$ satisfies $\alpha\leq A(2R)$ in $B(p,2R)$. Moreover, the non-Riemannian tensors satisfy $F(U)+F^{*}(\mathcal{T})+F(\mathrm{div}C(V))\leq K_{0}$. Let $u$ be a bounded positive solution to the Finslerian Allen-Cahn equation (3) in $B(p,2R)$ with $u\leq C$ where $C$ is a positive constant.
(1) If $C\leq 1$, then we have
@@
in $B(p,R)$ for any $10<\epsilon<$ 1, where $C_{1},C_{2}$ are positive constants.
(2) If $C>1$, then we have
@@
Let $\tilde{\phi}\in C^{2}[0,+\infty)$ such that $\tilde{\phi}(r)=1$ when $r\leq 1$ and $\tilde{\phi}(r)=0$ whenever $r\geq 2$. Hence $0\leq\tilde{\phi}(r)\leq 1$. Moreover, we may assume that
@@
$\Delta^{\nabla u}(\phi H)=H\Delta^{\nabla u}\phi+\phi\Delta^{\nabla u}H+\frac{2}{\phi}g_{\nabla u}(\nabla^{\nabla u}(\phi H),\nabla^{\nabla u}\phi)-\frac{2H}{\phi}F_{\nabla u}^{2}(\nabla^{\nabla u}\phi),$ (63)
where $\nabla^{\nabla u}\phi=\tilde{\phi}^{\prime}\frac{\nabla^{\nabla u}r}{R}$, $\Delta^{\nabla u}\phi=\tilde{\phi}^{\prime\prime}\frac{F^{2}_{\nabla u}(\nabla^{\nabla u}r)}{R^{2}}+\tilde{\phi}^{\prime}\frac{\Delta^{\nabla u}r}{R}$. Plugging these two terms into (63) gives that
$\begin{split}\phi\Delta^{\nabla u}(\phi H)=&\phi H(\Delta^{\nabla u}\phi-\frac{2}{\phi}F_{\nabla u}^{2}(\nabla^{\nabla u}\phi))\\ &+2\phi g_{\nabla u}(\nabla^{\nabla u}(\phi H),\nabla^{\nabla u}\log\phi)+\phi^{2}\Delta^{\nabla u}H\\ \geq&\phi H\left[-\frac{\alpha C_{2}}{R^{2}}-\frac{2\alpha C_{1}^{2}}{R^{2}}-\frac{C_{1}}{R}(C(N,\alpha)\mathfrak{ct}_{l}(r)+\sqrt{\alpha}K_{0})\right]\\ &+2\phi g_{\nabla u}(\nabla^{\nabla u}(\phi H),\nabla^{\nabla u}\log\phi)+\phi^{2}\Delta^{\nabla u}H,\end{split}$ (64)
on $M_{u}$. (64) provides by taking the estimation of $\Delta^{\nabla u}H$ in Lemma 9 that
$\begin{split}\phi\Delta^{\nabla u}(\phi H)\geq&2\phi g_{\nabla u}(\nabla^{\nabla u}(\phi H),\nabla^{\nabla u}\log\phi)+\frac{2\phi}{q}g_{\nabla u}(\nabla^{\nabla u}(\phi H),\nabla^{\nabla u}\log w)\\ &-B\phi H+\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{(\phi H)^{2}}{s^{2}q^{2}}+2\phi^{2}\frac{Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\\ &+\phi^{2}\left[\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{(sq+s-1)^{2}}{s^{2}q^{2}}-2(\frac{1}{\epsilon}-1)\right]\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+\phi\frac{4(1-\epsilon)(sq+s-1)}{N-\epsilon(N-n)s^{2}q^{2}}(\phi H)\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}+2w^{-\frac{2}{q}}\phi^{2}H\\ &+(2-6s)\phi^{2}w^{-\frac{2}{q}}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}-\frac{2}{q}(\phi H)g_{\nabla u}(\nabla^{\nabla u}\phi,\nabla^{\nabla u}\log w),\end{split}$ (65)
where $B=\frac{2\alpha C_{1}^{2}}{R^{2}}+\frac{C_{1}}{R}\left[\sqrt{K(2R)C(N,A)}\coth\left(R\sqrt{\frac{K(2R)}{C(N,A)}}\right)+C_{0}(K_{0},A)\right]+\frac{\alpha C_{2}}{R^{2}}$. It could be deduced from (65) that
@@
$\begin{split}B\phi H\geq&\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{(\phi H)^{2}}{s^{2}q^{2}}+2\phi^{2}\frac{Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\\ &+\phi^{2}\left[\frac{2(1-\epsilon)}{N-\epsilon(N-n)}\frac{(sq+s-1)^{2}}{s^{2}q^{2}}-2(\frac{1}{\epsilon}-1)\right]\frac{F_{\nabla u}^{4}(\nabla^{\nabla u}w)}{w^{4}}\\ &+\phi\frac{4(1-\epsilon)(sq+s-1)}{N-\epsilon(N-n)s^{2}q^{2}}(\phi H)\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}\\ &+(2-6s)\phi^{2}w^{-\frac{2}{q}}\frac{F_{\nabla u}^{2}(\nabla^{\nabla u}w)}{w^{2}}-\frac{2}{q}(\phi H)g_{\nabla u}(\nabla^{\nabla u}\phi,\nabla^{\nabla u}\log w),\end{split}$ (67)
is valid at $x_{1}$. Next we also consider (67) in two cases: (1) $C\leq 1$ and (2) $C>1$.
@@
$\frac{2Ric^{N}(\nabla^{\nabla u}w)}{w^{2}}\phi^{2}\geq-2K(2R)\phi^{2}\frac{F^{2}(\nabla^{\nabla u}w)}{w^{2}}\geq-2AK(2R)\phi H,$ (68)